# On the Similarities of Embeddings in Contrastive Learning

### v1

@@ Abstract
Contrastive learning (CL) operates on a simple yet effective principle: embeddings of positive pairs are pulled together, while those of negative pairs are pushed apart. Although various forms of contrastive loss have been proposed and analyzed from different perspectives, prior works lack a comprehensive framework that systematically explains a broad class of these objectives. In this paper, we present a unified framework for understanding CL, which is based on analyzing the cosine similarity between embeddings of positive and negative pairs. In full-batch settings, we show that perfect alignment of positive pairs is unattainable when similarities of negative pairs fall below a certain threshold, and that this misalignment can be alleviated by incorporating within-view negative pairs. In mini-batch settings, we demonstrate that smaller batch sizes incur stronger separation among negative pairs within batches, which leads to higher variance in similarities of negative pairs. To address this limitation of mini-batch CL, we introduce an auxiliary loss term that reduces the variance of similarities of negative pairs in CL. Empirical results demonstrate that incorporating the proposed loss consistently improves the performance of CL methods in small-batch training.
@@ 1 Introduction
Contrastive learning (CL) has emerged as a powerful approach in representation learning (Chen & He, 2021; Chen et al., 2020; He et al., 2020; Radford et al., 2021; Zhai et al., 2023; Khosla et al., 2020). In CL, an embedding model is trained to produce similar representations for two different views of the same data instance—referred to as a positive pair—and dissimilar representations for views from different data instances—referred to as negative pairs. Recent empirical studies have shown that representations learned by these objectives, i.e., aligning positive pairs while separating negative pairs, achieve remarkable performance across various downstream tasks (Wang & Isola, 2020).
@@
In parallel, several theoretical studies have analyzed the embeddings learned by CL in the full-batch training setting. Lu & Steinerberger (2022) showed that the widely used InfoNCE loss (Oord et al., 2018) achieves its minimum value when positive pairs are perfectly aligned, and negative pairs are uniformly separated with the cosine similarity of $-\tfrac{1}{n-1}$, where $n$ is the size of the training dataset. Lee et al. (2024) further extended this optimality analysis to other contrastive loss functions, including the recently proposed SigLIP loss (Zhai et al., 2023).
@@
However, due to computational constraints, CL methods are typically implemented using mini-batches rather than full batches in practical scenarios. This has motivated recent studies on understanding whether the property of the optimal state observed in full-batch training, i.e., perfect alignment of positive pairs and uniform separation of negative pairs with the similarity of $-\tfrac{1}{n-1}$, also holds under the mini-batch setting. Cho et al. (2024) partially addressed this by showing that, for the InfoNCE loss, the optimal embeddings of mini-batch training are identical to those of full-batch training, only when the sum of all possible mini-batch losses is minimized. Further analysis by Koromilas et al. (2024) explored embeddings learned through kernel-based contrastive losses (Li et al., 2021; Waida et al., 2023) in mini-batch settings. While these studies provide valuable insights into understanding CL, their analyses are limited to specific classes of contrastive loss functions and do not readily generalize to broader forms of CL objectives.
@@
To address this limitation, we propose a unified framework for analyzing the embeddings learned by both full-batch and mini-batch CL. Our analysis centers on the cosine similarity between embeddings of both positive and negative pairs. By characterizing the statistical properties of similarities of these embeddings, we reveal how different contrastive loss functions influence the structure of the learned embedding space under various training conditions. Our key contributions are summarized as follows:
@@
  • In full-batch CL, we identify a fundamental trade-off between the alignment of positive pairs and the separation of negative pairs. Specifically, the perfect alignment of positive pairs is not feasible when the average similarity of negative pairs falls below the threshold $-\tfrac{1}{n-1}$. We show that such misalignment arises in a class of existing contrastive losses, which can be mitigated by incorporating within-view negative pairs into the loss.
@@
  • In mini-batch CL, we demonstrate that negative pairs sampled within the same batch exhibit stronger separation compared to those drawn from different batches. As a result, we show that smaller batch sizes induce a higher variance in the similarities of negative pairs in the embedding space.
@@
  • To address the high variance of the similarities of negative pairs in mini-batch training, we introduce an auxiliary loss term that can be integrated into contrastive loss functions to explicitly reduce this variance. Empirical results show that incorporating the proposed term improves the performance of CL methods, especially in small-batch training scenarios.
@@ 2 Related Work
The InfoNCE loss (Gutmann & Hyvärinen, 2010; Oord et al., 2018) is a widely adopted contrastive loss function that has been applied to various tasks (Wu et al., 2018; Hjelm et al., 2019; Bachman et al., 2019; Chi et al., 2021; Gao et al., 2021; Qian et al., 2021). SimCLR (Chen et al., 2020) modifies the InfoNCE loss to improve robustness to augmentations by treating different augmented views of the same instance as positives and all other augmented instances in the batch as negatives. However, SimCLR simultaneously optimizes both positive and negative pairs in the normalization, which can introduce conflicts during optimization. To address this issue, Decoupled Contrastive Loss (DCL) (Yeh et al., 2022) modifies the normalization so that the selection of negative pairs is restricted. Building on DCL, Decoupled Hyperspherical Energy Loss (DHEL) (Koromilas et al., 2024) further refines the selection by focusing on negative pairs between augmented views of the same instance, which are more challenging than those between different instances. Figure 5 visualizes which pairs are included as positives and negatives for each method.
@@
Several studies have examined how optimal embeddings should be structured to minimize contrastive loss (Lee et al., 2025). Lu & Steinerberger (2022) showed that the optimal embeddings that minimize the InfoNCE loss form a simplex Equiangular Tight Frame (ETF) (Papyan et al., 2020; Sustik et al., 2007), where each positive pair is perfectly aligned and negative pairs are equally separated at the same angle, resulting in maximal separation among representations. Lee et al. (2024) further showed that the sigmoid-based contrastive loss, a variant of the softmax-based InfoNCE loss, also achieves the same simplex ETF optimum when the temperature parameter of the loss is sufficiently large. This optimal simplex ETF structure still holds even in mini-batch training, provided that optimization is performed over all possible mini-batch combinations, rather than a single batch at a time (Cho et al., 2024). Building on these findings, we introduce a unified theoretical analysis of the similarities between embedding pairs.
@@
CL shows outstanding performance, particularly when trained with large batch sizes (Chen et al., 2020; Radford et al., 2021; Pham et al., 2023; Tian et al., 2020b; Jia et al., 2021). However, large batch sizes require substantial memory resources, which poses practical challenges and often necessitates the use of smaller batches. This compromise in batch size typically leads to performance degradation, motivating several theoretical studies to investigate its causes (Cho et al., 2024; Koromilas et al., 2024). For example, Yuan et al. (2022) demonstrate that the optimization error in SimCLR (Chen et al., 2020) is upper bounded by a function inversely proportional to the batch size, indicating that smaller batches yield larger optimization errors. Additionally, Chen et al. (2022) show that contrastive loss functions exhibit increasing discrepancies between the true gradients and those estimated during training as the batch size decreases. While previous studies have primarily focused on optimization error and gradient estimation, we prove that training with small batch sizes leads to increased variance in the similarities of negative pairs in learned embeddings.
@@ 3 Problem Setup
The encoder produces outputs ${\bm{u}}=f({\bm{x}})$ and ${\bm{v}}=f({\bm{y}})$, which together form an embedding pair $({\bm{u}},{\bm{v}})$. When the embedding pair is generated from augmented views of the same instance, it is called a positive embedding pair and is encouraged to be similar. In contrast, a negative embedding pair, where each embedding comes from different instances, is encouraged to be dissimilar. For simplicity, we refer to positive embedding pairs as positive pairs and similarly to negative pairs when there is no risk of confusion.
@@
Figure 1: Illustration of negative pair considered in the loss formulations defined in Def. 3.1 and Def. 3.2, which depends on the choice of $(c_{1},c_{2})$. Each grid shows all possible pairs between embeddings in ${\bm{U}}_{[n]}$ and ${\bm{V}}_{[n]}$, and each cell represents one pair. Green regions represent positive pairs, and blue-striped regions indicate which negative pairs are included in the loss.
@@
For any $a\leq b$ with $a,b\in[n]$, let the set of embedding pairs be denoted by $({\bm{U}}_{[a:b]},{\bm{V}}_{[a:b]}):=\{({\bm{u}}_{i},{\bm{v}}_{i}):i\in[a:b]\}$, derived from instances whose indices range from $a$ to $b$. The notation $({\bm{U}}_{[a:b]},{\bm{V}}_{[a:b]})\sim f_{\sharp}\hat{p}_{\operatorname{pos}}^{[a:b]}$ indicates that each positive pair $({\bm{u}}_{i},{\bm{v}}_{i})$ is sampled from $f_{\sharp}\hat{p}_{\operatorname{pos}}^{i}$ for all $i\in[a:b]$. Note that ${\bm{u}}_{i}$ and ${\bm{v}}_{i}$ are random variables, as they are the encoder outputs of randomly augmented views. For simplicity, the subscript is omitted for the set of all $n$ pairs, i.e., $({\bm{U}},{\bm{V}}):=({\bm{U}}_{[n]},{\bm{V}}_{[n]})$.
@@
where $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ denotes the contrastive loss function for a given sample $({\bm{U}},{\bm{V}})$. In this work, we focus on two specific forms of contrastive loss functions used in practice.
@@
$\mathcal{L}_{\operatorname{ind-add}}({\bm{U}}_{{\bm{I}}},$ ${\bm{V}}_{{\bm{I}}}):=-\frac{1}{n}\sum_{i\in{\bm{I}}}\phi({\bm{u}}_{i}^{\top}{\bm{v}}_{i})$ (3)
$+\frac{c_{1}}{n(n-1)}\sum_{i\neq j\in{\bm{I}}}\psi({\bm{u}}_{i}^{\top}{\bm{v}}_{j})$
$+\frac{c_{2}}{2n(n-1)}\sum_{i\neq j\in{\bm{I}}}\left(\psi({\bm{u}}_{i}^{\top}{\bm{u}}_{j})+\psi({\bm{v}}_{i}^{\top}{\bm{v}}_{j})\right)$
for some constants $(c_{1},c_{2})\in\{(0,1),(1,0),(1,1)\}$, where $\phi:{\mathbb{R}}\to{\mathbb{R}}$ is a differentiable, concave, and increasing function, and $\psi:{\mathbb{R}}\to{\mathbb{R}}$ is a differentiable, convex, and increasing function. Here, $i\neq j\in[n]$ is a simplified notation representing $i\in[n]$ and $j\in[n]\setminus\{i\}$.
@@
The first form of losses in Def. 3.1 encompasses a variety of contrastive losses such as InfoNCE (Oord et al., 2018; Radford et al., 2021), SimCLR (Chen et al., 2020), DCL (Yeh et al., 2022), and DHEL (Koromilas et al., 2024), see Appendix A.1. The second form in Def. 3.2 includes CL losses such as SigLIP (Zhai et al., 2023) and Spectral CL (HaoChen et al., 2021), see Appendix A.2.
The difference between the two forms of loss functions lies in computational efficiency. The first form in Def. 3.1 necessitates simultaneous computation based on pairwise similarities across the entire set of embeddings due to the need for normalization, which becomes impractical for extremely large batch sizes. In contrast, the second form in Def. 3.2 is independently additive, allowing it to compute components of each pairwise similarity individually and aggregate them, making it applicable to larger datasets.
@@
Figure 2: Visualization of three different cases of eight embeddings on the three-dimensional unit sphere. Positive embedding pairs are represented in the same color and share the same subscript, while negative pairs refer to any two embeddings with different subscripts. In (a) and (b), embeddings minimize the fixed mini-batch contrastive loss described in Theorem 5.5, with the batches partitioned as $\{{\bm{u}}_{1},{\bm{v}}_{1},{\bm{u}}_{2},{\bm{v}}_{2}\}$ and $\{{\bm{u}}_{3},{\bm{v}}_{3},{\bm{u}}_{4},{\bm{v}}_{4}\}$. In (c), the embeddings minimize the full-batch contrastive loss described in Theorem 5.1. In all cases, the expectation of negative-pair similarities, $\mathbb{E}_{i\neq j\in[4]}[{\bm{u}}_{i}^{\top}{\bm{v}}_{j}]$, remains the same. However, the variance of negative-pair similarities, $\mathrm{Var}_{i\neq j\in[4]}[{\bm{u}}_{i}^{\top}{\bm{v}}_{j}]$, increases in mini-batch trainings, indicating that some negative pairs are much more similar to each other while others are more dissimilar.

### v2

@@ Abstract
Contrastive learning operates on a simple yet effective principle: Embeddings of positive pairs are pulled together, while those of negative pairs are pushed apart. In this paper, we propose a unified framework for understanding contrastive learning through the lens of cosine similarity, and present two key theoretical insights derived from this framework. First, in full-batch settings, we show that perfect alignment of positive pairs is unattainable when negative-pair similarities fall below a threshold, and this misalignment can be mitigated by incorporating within-view negative pairs into the objective. Second, in mini-batch settings, smaller batch sizes induce stronger separation among negative pairs in the embedding space, i.e., higher variance in their similarities, which in turn degrades the quality of learned representations compared to full-batch settings. To address this, we propose an auxiliary loss that reduces the variance of negative-pair similarities in mini-batch settings. Empirical results show that incorporating the proposed loss improves performance in small-batch settings.
@@ 1 Introduction
Contrastive learning (CL) has emerged as a powerful approach to representation learning (Chen & He, 2021; Chen et al., 2020; He et al., 2020; Radford et al., 2021; Zhai et al., 2023; Khosla et al., 2020). In CL, an embedding model is trained to produce similar representations for two different views of the same data instance—referred to as a positive pair—and dissimilar representations for views from different data instances—referred to as negative pairs. Recent empirical studies have shown that representations learned by these objectives, i.e., aligning positive pairs while separating negative pairs, achieve remarkable performance across various downstream tasks (Wang & Isola, 2020).
@@
In parallel, several theoretical studies have analyzed the embeddings obtained by CL in full-batch settings. Lu & Steinerberger (2022) showed that the widely used InfoNCE loss (Oord et al., 2018) achieves its minimum value when positive pairs are perfectly aligned, and negative pairs are uniformly separated with the cosine similarity of $-\tfrac{1}{n-1}$, where $n$ is the size of the training dataset. Lee et al. (2024) further extended this optimality analysis to other contrastive losses, including the SigLIP loss (Zhai et al., 2023).
@@
Due to computational constraints, CL methods are typically implemented using mini-batches rather than full batches in practical scenarios. This has motivated recent studies on understanding whether the property of the optimal state observed in full-batch settings, i.e., perfect alignment of positive pairs and uniform separation of negative pairs with the similarity of $-\tfrac{1}{n-1}$, also holds under mini-batch settings. Cho et al. (2024) partially addressed this by showing that, for the InfoNCE loss, the optimal embeddings of mini-batch settings are identical to those of full-batch settings, only when the sum of all possible mini-batch losses is minimized. Koromilas et al. (2024) explored embeddings learned through kernel-based contrastive losses (Li et al., 2021; Waida et al., 2023) in mini-batch settings. While these studies provide valuable insights into understanding CL, their analyses are limited to specific forms of contrastive losses.
@@
To address this limitation, we propose a unified framework for analyzing the embeddings obtained by both full-batch and mini-batch settings. Our analysis centers on the cosine similarity of embeddings of both positive and negative pairs. By characterizing the statistical properties of similarities of these embeddings, we reveal how different contrastive losses influence the structure of the learned embedding space under various training conditions.
Our key contributions are summarized as follows:
@@
  • In full-batch settings, we identify a fundamental trade-off between the alignment of positive pairs and the separation of negative pairs. Specifically, we show that the perfect alignment of positive pairs is not feasible when the average similarity of negative pairs falls below the certain threshold $-\tfrac{1}{n-1}$. We demonstrate that such misalignment arises in a class of existing contrastive losses, which can be mitigated by incorporating within-view negative pairs into the contrastive loss.
@@
  • In mini-batch settings, we demonstrate that negative pairs within the same batch exhibit stronger separation compared to those from different batches. As a result, we show that smaller batch sizes induce a higher variance in the similarities of negative pairs in the embedding space. We identify this increased variance as a distinctive feature of mini-batch settings that is absent in full-batch settings.
@@
  • Motivated by prior studies that full-batch settings often outperform their mini-batch counterparts, we hypothesize that the increased variance may underlie this performance gap. To explore this, we propose an auxiliary loss that can be integrated into contrastive losses to reduce this variance. Empirical results show that incorporating the proposed term improves the performance of CL methods, especially in small-batch settings.
@@ 2 Related Work
The InfoNCE loss (Gutmann & Hyvärinen, 2010; Oord et al., 2018) is a widely adopted contrastive loss that has been applied to various tasks (Wu et al., 2018; Hjelm et al., 2019; Bachman et al., 2019; Chi et al., 2021; Gao et al., 2021; Qian et al., 2021). SimCLR (Chen et al., 2020) modifies the InfoNCE loss to improve robustness to augmentations by treating different augmented views of the same instance as positives and all other augmented instances in the batch as negatives. However, SimCLR simultaneously optimizes both positive and negative pairs in the normalization, which can introduce conflicts during optimization. To address this issue, Decoupled Contrastive Loss (DCL) (Yeh et al., 2022) modifies the normalization so that the selection of negative pairs is restricted. Building on DCL, Decoupled Hyperspherical Energy Loss (DHEL) (Koromilas et al., 2024) further refines the selection by focusing on negative pairs between augmented views of the same instance, which are more challenging than those between different instances. Figure 5 visualizes which pairs are included as positives and negatives for each method.
@@
Several studies have examined how optimal embeddings should be structured to minimize contrastive loss (Lee et al., 2025). Lu & Steinerberger (2022) showed that the optimal embeddings that minimize the InfoNCE loss form a simplex Equiangular Tight Frame (ETF) (Papyan et al., 2020; Sustik et al., 2007), where each positive pair is perfectly aligned and negative pairs are equally separated at the same angle, resulting in maximal separation among representations. Lee et al. (2024) further showed that the sigmoid-based contrastive loss, a variant of the softmax-based InfoNCE loss, also achieves the same simplex ETF optimum when the temperature parameter of the loss is sufficiently large. This optimal simplex ETF structure still holds even in mini-batch settings, provided that optimization is performed over all possible mini-batch combinations, rather than a single batch at a time (Cho et al., 2024). Building on these findings, we introduce a unified theoretical analysis of the similarities of embedding pairs.
@@
CL shows outstanding performance, particularly when trained with large batch sizes (Chen et al., 2020; Radford et al., 2021; Pham et al., 2023; Tian et al., 2020b; Jia et al., 2021). However, large batch sizes require substantial memory resources, which poses practical challenges and often necessitates the use of smaller batches. This compromise in batch size typically leads to performance degradation, motivating several theoretical studies to investigate its causes (Cho et al., 2024; Koromilas et al., 2024). For example, Yuan et al. (2022) demonstrate that the optimization error in SimCLR (Chen et al., 2020) is upper bounded by a function inversely proportional to the batch size, indicating that smaller batches yield larger optimization errors. Additionally, Chen et al. (2022) show that contrastive losses exhibit increasing discrepancies between the true gradients and those estimated during training as the batch size decreases. While previous studies have primarily focused on optimization error and gradient estimation, we prove that training with small batch sizes leads to increased variance in the similarities of negative pairs in learned embeddings.
@@ 3 Problem Setup
The encoder produces outputs ${\bm{u}}=f({\bm{x}})$ and ${\bm{v}}=f({\bm{y}})$, which together form an embedding pair $({\bm{u}},{\bm{v}})$. When the embedding pair is generated from augmented views of the same instance, it is called a positive embedding pair and is encouraged to be similar. In contrast, a negative embedding pair, where each embedding comes from different instances, is encouraged to be dissimilar. For simplicity, we refer to positive embedding pairs as positive pairs and similarly to negative pairs, when there is no risk of confusion.
@@
Figure 1: Illustration of negative pair considered in the loss formulations defined in Def. 3.1 and Def. 3.2, which depends on the choice of $(c_{1},c_{2})$. Each grid shows all possible pairs of embeddings in ${\bm{U}}_{[n]}$ and ${\bm{V}}_{[n]}$, and each cell represents one pair. Green regions represent positive pairs, and blue-striped regions indicate which negative pairs are included in the loss.
@@
For any $a\leq b$ with $a,b\in[n]$, let $({\bm{U}}_{[a:b]},{\bm{V}}_{[a:b]}):=\{({\bm{u}}_{i},{\bm{v}}_{i}):i\in[a:b]\}$ be the set of embedding pairs for instances whose indices range from $a$ to $b$. The notation $({\bm{U}}_{[a:b]},{\bm{V}}_{[a:b]})\sim f_{\sharp}\hat{p}_{\operatorname{pos}}^{[a:b]}$ indicates that each positive pair $({\bm{u}}_{i},{\bm{v}}_{i})$ is sampled from $f_{\sharp}\hat{p}_{\operatorname{pos}}^{i}$ for all $i\in[a:b]$. Note that ${\bm{u}}_{i}$ and ${\bm{v}}_{i}$ are random variables, as they are the encoder outputs of randomly augmented views. For simplicity, the subscript is omitted for the set of all $n$ pairs, i.e., $({\bm{U}},{\bm{V}}):=({\bm{U}}_{[n]},{\bm{V}}_{[n]})$.
@@
where $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ denotes the contrastive loss for a given sample $({\bm{U}},{\bm{V}})$. In this work, we focus on two specific forms of contrastive losses used in practice:
@@
$\mathcal{L}_{\operatorname{ind-add}}({\bm{U}}_{{\bm{I}}}$ $,{\bm{V}}_{{\bm{I}}}):=-\frac{1}{|{\bm{I}}|}\sum_{i\in{\bm{I}}}\phi({\bm{u}}_{i}^{\top}{\bm{v}}_{i})$ (3)
$+\frac{c_{1}}{|{\bm{I}}|(|{\bm{I}}|-1)}\sum_{i\neq j\in{\bm{I}}}\psi({\bm{u}}_{i}^{\top}{\bm{v}}_{j})$
$+\frac{c_{2}}{2|{\bm{I}}|(|{\bm{I}}|-1)}\sum_{i\neq j\in{\bm{I}}}\!\!\left(\psi({\bm{u}}_{i}^{\top}{\bm{u}}_{j})+\psi({\bm{v}}_{i}^{\top}{\bm{v}}_{j})\right)$
for some constants $(c_{1},c_{2})\in\{(0,1),(1,0),(1,1)\}$, where $\phi:{\mathbb{R}}\to{\mathbb{R}}$ is a differentiable, concave, and increasing function, and $\psi:{\mathbb{R}}\to{\mathbb{R}}$ is a differentiable, convex, and increasing function. Here, $i\neq j\in{\bm{I}}$ is a simplified notation representing $i\in{\bm{I}}$ and $j\in{\bm{I}}\setminus\{i\}$.
@@
The first form of losses in Def. 3.1 encompasses a variety of contrastive losses such as InfoNCE (Oord et al., 2018; Radford et al., 2021), SimCLR (Chen et al., 2020), DCL (Yeh et al., 2022), and DHEL (Koromilas et al., 2024), see Appendix A.1. The second form in Def. 3.2 includes contrastive losses such as SigLIP (Zhai et al., 2023) and Spectral CL (HaoChen et al., 2021), see Appendix A.2.
The difference between the two forms of losses lies in computational efficiency. The first form in Def. 3.1 necessitates simultaneous computation based on pairwise similarities across the entire set of embeddings due to the need for normalization, which becomes impractical for extremely large batch sizes. In contrast, the second form in Def. 3.2 is independently additive, allowing it to compute components of each pairwise similarity individually and aggregate them, making it applicable to larger datasets.
@@
Figure 2: Visualization of three different cases of eight embeddings on the three-dimensional unit sphere. Positive embedding pairs are represented in the same color and share the same subscript, while negative pairs refer to any two embeddings with different subscripts. In (a) and (b), embeddings minimize the fixed mini-batch contrastive loss described in Theorem 5.5, with the batches partitioned as $\{{\bm{u}}_{1},{\bm{v}}_{1},{\bm{u}}_{2},{\bm{v}}_{2}\}$ and $\{{\bm{u}}_{3},{\bm{v}}_{3},{\bm{u}}_{4},{\bm{v}}_{4}\}$. In (c), the embeddings minimize the full-batch contrastive loss described in Theorem 5.1. In all cases, the expectation of negative-pair similarities, $\mathbb{E}_{i\neq j\in[4]}[{\bm{u}}_{i}^{\top}{\bm{v}}_{j}]$, remains the same. However, the variance of negative-pair similarities, $\mathrm{Var}_{i\neq j\in[4]}[{\bm{u}}_{i}^{\top}{\bm{v}}_{j}]$, increases in mini-batch settings, indicating that some negative pairs are much more similar to each other while others are more dissimilar.